The following selected data are adapted from the November 30, 2014, and November 24, 2013, consolidated balance sheets and income statements for the years then ended for Levi Strauss & Co. and Subsidiaries. All amounts are reported in thousands. Required: Calculate the missing amounts for each year. (Enter your answers in thousands of dollars.)

                                                                                                     2014                                   2013

                                                     Net revenues                      4099147        

                                                     Cost of goods sold                                                           2516521       

                                                     Gross profit                         1782190                               2057230

                 Selling general  and administrative expenses                                                     1,276,530

                                            Operating Income                         

              Interest expense and other expenses, net               268,529                                 344,641

                                   Income before income taxes                 91610                                                   

                                       Income tax expense                                                                      318,025

                                      Net Income                                       26475                                   118034
Answer #1

2014

2013

Net revenues

$ 40,99,147.00

$ 45,73,751.00

Cost of goods sold

$ 23,16,957.00

$ 25,16,521.00

Gross profit

$ 17,82,190.00

$ 20,57,230.00

Selling and administrative expenses

$ 14,22,051.00

$ 12,76,530.00

Operating income

$     3,60,139.00

$    7,80,700.00

Interest expenses and other expenses, net

$     2,68,529.00

$    3,44,641.00

Income before income taxes

$        91,610.00

$    4,36,059.00

Income tax expense

$        65,135.00

$    3,18,025.00

Net income

$        26,475.00

$    1,18,034.00

Working

2014

2013

Net revenues

4099147

=2516521+2057230

Cost of goods sold

=4099147-1782190

2516521

Gross profit

1782190

2057230

Selling and administrative expenses

=1782190-360139

1276530

Operating income

=268529+91610

=2057230-1276530

Interest expenses and other expenses, net

268529

344641

Income before income taxes

91610

=780700-344641

Income tax expense

=91610-26475

318025

Net income

26475

118034